# Hybrid Feedback for Three-dimensional Convex Obstacle Avoidance (Extended version)

### v4

@@ Abstract
We propose a hybrid feedback control scheme for the autonomous robot navigation problem in three-dimensional environments with arbitrarily-shaped convex obstacles. The proposed hybrid control strategy, which consists in switching between the move-to-target mode and the obstacle-avoidance mode, guarantees global asymptotic stability of the target location in the obstacle-free workspace. We also provide a procedure for the implementation of the proposed hybrid controller in a priori unknown environments and validate its effectiveness through simulation results.
@@ I INTRODUCTION
Safe autonomous robot navigation consists in steering a robot to a target location while avoiding obstacles. One common method is the Navigation Function (NF) approach [1], where a combination of attractive and repulsive vector fields, with a properly tuned parameter, guarantees safe, almost [^1] Almost global convergence refers to convergence from all initial locations except a set of zero Lebesgue measure. global convergence of the robot to the target location in sphere worlds. For general convex and star-shaped obstacles, diffeomorphic mappings [1, 2] allow the use of the NF approach but require complete knowledge of the environment. In [3], the NF method has been extended to convex obstacles that satisfy certain curvature properties, and are not too flat and not too close to the target. The work in [4] removes this flatness limitation, in environments with ellipsoidal obstacles, by using the Hessian to locally transform areas near obstacles into spherical regions. However, similar to [3], it is assumed that the entire shape of the obstacle is known when the robot visits its neighborhood.
@@
In [5, 6, 7], the authors combine control Lyapunov functions (CLF) and control barrier functions (CBF) to design feedback laws that ensure safe convergence to a target location. The authors in [6] introduced a CBF-based method for multi-robot navigation in two-dimensional environments with circular obstacles.A comparative analysis between the APF and CBF approaches can be found in [7]. As shown in [8] and [9], CLF-CBF methods, like the NF approach, suffer from the undesired equilibria problem and guarantee, at best, almost global asymptotic stability in sphere worlds.
In [10], a feedback controller based on Nagumo’s theorem [11, Theorem 4.7] was proposed, guaranteeing safe, almost global convergence to the target location in a priori unknown environments with strongly convex obstacles. The work in [12] introduced a reactive navigation approach using separating hyperplanes for robots in environments with unknown, sufficiently separated, strongly convex obstacles.
@@
However, the approaches discussed above provide, at best, almost global convergence guarantees due to the undesired equilibria that are generated when using continuous time-invariant vector fields [13]. This can be resolved by introducing discontinuities in the control, as shown in [14, 15, 16, 17].
@@
In [14], a discontinuous feedback law was developed for robot navigation in partially known two-dimensional environments, guiding the robot with the navigation function gradient near known obstacles and boundary-following near unknown ones. In [15], the authors proposed a discontinuous feedback controller for autonomous navigation of nonholonomic robots in two-dimensional environments with non-convex obstacles, imposing restrictions on the curvatures of obstacles’ boundaries and the inter-obstacle arrangements. In [16], the authors proposed a hybrid control law to globally asymptotically stabilize a class of linear systems with drift while avoiding neighborhoods of unsafe isolated points. In [17], hybrid control techniques were employed to achieve global stabilization of target locations in $n$-dimensional environments with sufficiently separated ellipsoidal obstacles.
@@
In our earlier work [18, 19], we proposed a hybrid feedback control strategy for robot navigation in two-dimensional environments with convex and non-convex obstacles. In this paper, we extend this approach to ensure autonomous robot navigation in unknown $3$-dimensional environments with arbitrarily-shaped convex obstacles, including those with non-smooth boundaries. Our solution guarantees safe, global asymptotic convergence of the robot to the target location, unlike the methods in [4, 10, 12], which guarantee, at best, almost global asymptotic convergence. The approach in [17] applies to environments with ellipsoidal obstacles, the work in [10, 12] is limited to strongly convex obstacles, and in [14, 15] the obstacles are required to have smooth boundaries. In contrast, the proposed hybrid feedback controller works for arbitrarily-shaped convex obstacles with no restrictions on inter-obstacle arrangement, except for a mild separation condition. We also propose a sensor-based version of our approach for autonomous navigation in a priori unknown $3$-dimensional environments with arbitrarily-shaped convex obstacles.
@@ II-A Notations
The sets of real and natural numbers are denoted by $\mathbb{R}$ and $\mathbb{N}$, respectively. We identify vectors using bold lowercase letters. The Euclidean norm of a vector $\mathbf{p}\in\mathbb{R}^{n}$ is denoted by $\lVert\mathbf{p}\rVert$, and an Euclidean ball of radius $r\geq 0$ centered at $\mathbf{p}$ is represented by $\mathcal{B}_{r}(\mathbf{p})=\{\mathbf{q}\in\mathbb{R}^{n}|\lVert\mathbf{q}-\mathbf{p}\rVert\leq r\}.$ The set of $n-$dimensional unit vectors is given by $\mathbb{S}^{n-1}=\{\mathbf{p}\in\mathbb{R}^{n}|\lVert\mathbf{p}\rVert=1\}.$ The identity matrix of order $n$ is denoted by $\mathbf{I}_{n}$. The three-dimensional Special Orthogonal group is defined by $SO(3):=\{\mathbf{R}\in\mathbb{R}^{3\times 3}:\mathbf{R}^{\top}\mathbf{R}=\mathbf{I}_{3},\det(\mathbf{R})=1\}.$ For a given vector $\mathbf{x}:=[x_{1},x_{2},x_{3}]^{\top}\in\mathbb{R}^{3}$, we define $\mathbf{x}^{\times}$ as the skew-symmetric matrix, which is given by
$\mathbf{x}^{\times}=\begin{bmatrix}0&-x_{3}&x_{2}\\ x_{3}&0&-x_{1}\\ -x_{2}&x_{1}&0\end{bmatrix}.$ (1)
For two sets $\mathcal{A},\mathcal{B}\subset\mathbb{R}^{n}$, the relative complement of $\mathcal{B}$ with respect to $\mathcal{A}$ is denoted by $\mathcal{A}\setminus\mathcal{B}=\{\mathbf{a}\in\mathcal{A}|\mathbf{a}\notin\mathcal{B}\}$. The symbols $\partial\mathcal{A},\mathcal{A}^{\circ}$, $\mathcal{A}^{c}$ and $\bar{\mathcal{A}}$ represent the boundary, interior, complement and the closure of the set $\mathcal{A}$, respectively, where $\partial\mathcal{A}=\bar{\mathcal{A}}\backslash\mathcal{A}^{\circ}$. The cardinality of a set $\mathcal{A}$ is denoted by $\mathbf{card}(\mathcal{A})$. The Minkowski sum of the sets $\mathcal{A}$ and $\mathcal{B}$ is denoted by, $\mathcal{A}\oplus\mathcal{B}=\{\mathbf{a}+\mathbf{b}|\mathbf{a}\in\mathcal{A},\mathbf{b}\in\mathcal{B}\}$. The dilated version of a set $\mathcal{A}\subset\mathbb{R}^{n}$ with $r\geq 0$ is represented by $\mathcal{D}_{r}(\mathcal{A})=\mathcal{A}\oplus\mathcal{B}_{r}(\mathbf{0})$. The $r-$neighborhood of a set $\mathcal{A}$ is denoted by $\mathcal{N}_{r}(\mathcal{A})=\mathcal{D}_{r}(\mathcal{A})\setminus\mathcal{A}^{\circ}$ where $r$ is a strictly positive scalar.
@@ II-B Projection on a set
The set $\mathcal{PJ}(\mathbf{x},\mathcal{A})\subset\mathcal{A}$, which is defined as
$\mathcal{PJ}(\mathbf{x},\mathcal{A})=\{\mathbf{q}\in\mathcal{A}|\lVert\mathbf{x}-\mathbf{q}\rVert=d(\mathbf{x},\mathcal{A})\},$ (3)
is the set of points in $\mathcal{A}$ that are at the distance of $d(\mathbf{x},\mathcal{A})$ from $\mathbf{x}$. If $\mathbf{card}(\mathcal{PJ}(\mathbf{x},\mathcal{A}))$ is one, then the element of the set $\mathcal{PJ}(\mathbf{x},\mathcal{A})$ is denoted by $\Pi(\mathbf{x},\mathcal{A})$.
@@ II-C2 Line segment
The line segment joining two points $\mathbf{p}\in\mathbb{R}^{n}$ and $\mathbf{q}\in\mathbb{R}^{n}$ is given by
@@
II-C3 Hyperplane
@@
$\mathcal{P}(\mathbf{p},\mathbf{q}):=\{\mathbf{x}\in\mathbb{R}^{n}|\mathbf{q}^{\intercal}(\mathbf{x}-\mathbf{p})=0\}.$ (6)
The hyperplane divides the Euclidean space $\mathbb{R}^{n}$ into two half-spaces i.e., a closed positive half-space $\mathcal{P}_{\geq}(\mathbf{p},\mathbf{q})$ and a closed negative half-space $\mathcal{P}_{\leq}(\mathbf{p},\mathbf{q})$ which are obtained by substituting ‘$=$’ with $\geq$’ and $\leq$’ respectively, in the right-hand side of (6). We also use the notations $\mathcal{P}_{>}(\mathbf{p},\mathbf{q})$ and $\mathcal{P}_{<}(\mathbf{p},\mathbf{q})$ to denote the open positive and the open negative half-spaces such that $\mathcal{P}_{>}(\mathbf{p},\mathbf{q})=\mathcal{P}_{\geq}(\mathbf{p},\mathbf{q})\backslash\mathcal{P}(\mathbf{p},\mathbf{q})$ and $\mathcal{P}_{<}(\mathbf{p},\mathbf{q})=\mathcal{P}_{\leq}(\mathbf{p},\mathbf{q})\backslash\mathcal{P}(\mathbf{p},\mathbf{q})$.
II-D Hybrid system framework
@@
where the (set-valued) flow map $\mathbf{F}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ and jump map $\mathbf{J}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ govern continuous and discrete evolution, which can occur, respectively, in the flow set $\mathcal{F}\subset\mathbb{R}^{n}$ and the jump set $\mathcal{J}\subset\mathbb{R}^{n}$. The notions of solution $\phi$ to a hybrid system, hybrid time domain dom $\text{dom }\phi$, maximal and complete solution are, respectively, as in [20, Def. 2.6, Def. 2.3, Def. 2.7, p.30].

### v5

# N-dimensional Convex Obstacle Avoidance using Hybrid Feedback Control (Extended version)

@@ Abstract
This paper addresses the autonomous robot navigation problem in a priori unknown $n$-dimensional environments containing convex obstacles of arbitrary shapes and sizes. We propose a hybrid feedback control scheme that guarantees safe and global asymptotic convergence of the robot to a predefined target location. The proposed control strategy relies on a switching mechanism allowing the robot to operate either in the move-to-target mode or the obstacle-avoidance mode, based on its proximity to the obstacles and the availability of a clear straight path between the robot and the target. In the obstacle-avoidance mode, the robot is constrained to move within a two-dimensional plane that intersects the obstacle being avoided and the target, preventing it from retracing its path. The effectiveness of the proposed hybrid feedback controller is demonstrated through simulations in two-dimensional and three-dimensional environments.
@@ I Introduction
Safe autonomous robot navigation consists in steering a robot to a target location while avoiding obstacles. One commonly used navigation technique is the Artificial Potential Field (APF) approach [1], where a combination of attractive and repulsive vector fields guides the robot safely to the target location. However, this approach faces challenges with certain obstacle arrangements, leading to undesired stable local minima. The Navigation Function (NF) approach [2, 3] is effective in sphere world environments, addressing the local minima issue by limiting the repulsive field’s influence around the obstacles by means of a properly tuned parameter. However, this method ensures almost [^1] Almost global convergence refers to the convergence from all initial conditions except a set of zero Lebesgue measure. global convergence of the robot to the target location. To apply the NF approach to environments with general convex and star-shaped obstacles, one has to use diffeomorphic mappings from [2] and [4], which require global knowledge of the environment for implementation. In [5], the authors extended the NF approach to handle environments with convex obstacles with smooth boundaries that meet certain curvature conditions. However, this approach is limited to obstacles, which are not too flat and not too close to the target. In [6], for the case of ellipsoidal worlds, the authors removed the flatness limitation in [5], by providing a controller design, which locally transforms the region near the obstacle into a spherical region by using the Hessian information. However, similar to [5], it is assumed that the entire shape of the obstacle becomes known when the robot visits its neighbourhood.
@@
Other approaches, such as [7, 8, 9], rely on the control Lyapunov function (CLF) and the control barrier function (CBF) to design feedback control laws that achieve (simultaneously) convergence to a target set and avoidance of an unsafe set. In [8], the authors proposed a CBF-based method for multi-robot navigation in two-dimensional environments in the presence of circular obstacles. In [9], a comparative analysis between the APF-based and CBF-based approaches has been provided. However, the work in [10] and [11] demonstrates that the CLF-CBF-based navigation approach, similar to the NF approach, suffers from the undesired equilibria problem and provides (at best) almost global asymptotic stability guarantees in sphere worlds.
In [12], the authors proposed a feedback controller based on Nagumo’s theorem [13, Theorem 4.7] for autonomous navigation in environments with general convex obstacles. The forward invariance of the obstacle-free space is achieved by projecting the ideal velocity control vector, which points towards the target, onto the tangent cone at the obstacle boundary whenever it points towards the obstacle. This approach was extended in [14] to guarantee almost global asymptotic stability of the target location in a priori unknown environments containing strongly convex obstacles.
In [15], the authors proposed a purely reactive autonomous navigation approach based on separating hyperplanes for robots operating in environments cluttered with unknown but sufficiently separated convex obstacles with smooth boundaries which satisfy curvature conditions similar to [14]. This approach was extended in [16] to address partially known environments with non-convex obstacles, where the robot has geometric information about the non-convex obstacles but lacks precise knowledge of their locations in the workspace.
@@
The approaches discussed above provide, at best, almost global convergence guarantees due to the undesired equilibria that are generated when using continuous time-invariant vector fields [17]. This can be resolved by introducing discontinuities in the control, as shown in [18, 19, 20, 21, 22, 23, 24].
@@
In [18] and [19], hybrid control methods are employed to achieve robust global asymptotic stabilization in $\mathbb{R}^{2}$ for robots navigating towards a target location while avoiding collision with a single spherical obstacle. The approach in [18] has been extended in [20] to steer a group of planar robots in formation toward the source of an unknown but measurable signal while avoiding a single obstacle. In [21], the authors proposed a hybrid control law to globally asymptotically stabilize a class of linear systems with drift while avoiding neighbourhoods of unsafe isolated points. In [22], hybrid control techniques were employed to achieve global stabilization of target locations in $n$-dimensional environments with sufficiently separated ellipsoidal obstacles.
In [23], the authors proposed a discontinuous feedback-based autonomous navigation scheme for nonholonomic robots operating in two-dimensional environments with non-convex obstacles, subject to restrictions on inter-obstacle arrangements. In [24], a discontinuous feedback control law was introduced for autonomous robot navigation in partially known two-dimensional environments. When encountering a known obstacle, the control vector aligns with the negative gradient of the navigation function, whereas near unknown obstacles, the robot follows the boundary, using local curvature information.
@@
In our earlier works [25] and [26], we proposed a hybrid feedback control strategy for autonomous robot navigation in two-dimensional environments with arbitrarily-shaped convex and non-convex obstacles. In this paper, we extend this approach by proposing a hybrid control strategy with global asymptotic stability guarantees for autonomous navigation in a priori unknown $n$-dimensional environments containing arbitrarily-shaped convex obstacles, including those with non-smooth boundaries. The main contributions of the proposed research work are as follows:
  1. 1.
    Global asymptotic stability: The proposed autonomous navigation solution guarantees global asymptotic stabilization of the target location in unknown environments with convex obstacles of arbitrary shapes. Note that the few existing results in the literature achieving such strong stability results are of a hybrid type and are restricted to environments with ellipsoidal obstacles [22].
  2. 2.
    $n$-dimensional convex obstacles: The proposed hybrid feedback controller is applicable to $n$-dimensional workspaces containing convex obstacles of arbitrary shapes and sizes. In contrast, the autonomous navigation schemes in [23] and [24] are limited to two-dimensional environments, while the methods in [6] and [12] apply only to $n$-dimensional environments with ellipsoidal obstacles. The navigation approaches in [14] and [15] are restricted to environments with strongly convex obstacles.
  3. 3.
    Arbitrary interobstacle arrangements: There are no restrictions on the arrangement of obstacles, unlike those imposed in [23, Assumption 10] and [22, Theorem 2], except for the widely accepted mild condition in Assumption 9, which states that the robot can pass between any two obstacles while maintaining a positive distance.
  4. 4.
    Applicable in a priori unknown environments: Unlike the approaches in [5] and [6], which require global information about the obstacles, the proposed autonomous navigation method relies solely on range scanners and does not require a priori global knowledge of the environment (sensor-based technique).
The remainder of the paper is organized as follows. Section II introduces the notations and preliminaries used throughout the paper. The problem formulation is presented in Section III, followed by the proposed hybrid control algorithm in Section IV. Stability and safety guarantees of the navigation control scheme are discussed in Section V. In Section VI, the hybrid feedback control law is modified to ensure a monotonic decrease in distance to the target in sphere worlds. Section VII outlines the implementation procedure for the obstacle avoidance algorithm in a priori unknown environments. Simulation results are provided in Section VIII to demonstrate the algorithm’s effectiveness, and concluding remarks are given in Section IX.
@@ II-A Notations
The sets of real and natural numbers are denoted by $\mathbb{R}$ and $\mathbb{N}$, respectively. We identify vectors using bold lowercase letters. The Euclidean norm of a vector $\mathbf{p}\in\mathbb{R}^{n}$ is denoted by $\lVert\mathbf{p}\rVert$, and an Euclidean ball of radius $r\geq 0$ centered at $\mathbf{p}$ is represented by $\mathcal{B}_{r}(\mathbf{p})=\{\mathbf{q}\in\mathbb{R}^{n}|\lVert\mathbf{q}-\mathbf{p}\rVert\leq r\}.$ The set of $n-$dimensional unit vectors is given by $\mathbb{S}^{n-1}=\{\mathbf{p}\in\mathbb{R}^{n}|\lVert\mathbf{p}\rVert=1\}.$ The identity matrix of order $n$ is denoted by $\mathbf{I}_{n}$. The $n-$dimensional Special Orthogonal group is denoted by $SO(n):=\{\mathbf{R}\in\mathbb{R}^{n\times n}:\mathbf{R}^{\top}\mathbf{R}=\mathbf{I}_{n},\det(\mathbf{R})=1\}.$
For two sets $\mathcal{A},\mathcal{B}\subset\mathbb{R}^{n}$, the relative complement of $\mathcal{B}$ with respect to $\mathcal{A}$ is denoted by $\mathcal{A}\setminus\mathcal{B}=\{\mathbf{a}\in\mathcal{A}|\mathbf{a}\notin\mathcal{B}\}$. The symbols $\partial\mathcal{A},\mathcal{A}^{\circ}$, $\mathcal{A}^{c}$ and $\bar{\mathcal{A}}$ represent the boundary, interior, complement and the closure of the set $\mathcal{A}$, respectively, where $\partial\mathcal{A}=\bar{\mathcal{A}}\backslash\mathcal{A}^{\circ}$. The cardinality of a set $\mathcal{A}$ is denoted by $\mathbf{card}(\mathcal{A})$. The Minkowski sum of the sets $\mathcal{A}$ and $\mathcal{B}$ is denoted by $\mathcal{A}\oplus\mathcal{B}=\{\mathbf{a}+\mathbf{b}|\mathbf{a}\in\mathcal{A},\mathbf{b}\in\mathcal{B}\}$. The dilated version of a set $\mathcal{A}\subset\mathbb{R}^{n}$ with $r\geq 0$ is represented by $\mathcal{D}_{r}(\mathcal{A})=\mathcal{A}\oplus\mathcal{B}_{r}(\mathbf{0})$. The $r-$neighbourhood of a set $\mathcal{A}$ is denoted by $\mathcal{N}_{r}(\mathcal{A})=\mathcal{D}_{r}(\mathcal{A})\setminus\mathcal{A}^{\circ}$, where $r$ is a strictly positive scalar.
@@ II-B Projection on a set
If $\mathcal{A}$ is convex, the unique closest point to any $\mathbf{x}\in\mathbb{R}^{n}$ on $\mathcal{A}$ is denoted by $\Pi(\mathbf{x},\mathcal{A})$ and is defined as
$\Pi(\mathbf{x},\mathcal{O}_{i}):=\underset{\mathbf{q}\in\mathcal{A}}{\arg\min}\lVert\mathbf{x}-\mathbf{q}\rVert.$ (2)
@@ II-C2 Line segment
The line segment joining two points $\mathbf{p}\in\mathbb{R}^{n}$ and $\mathbf{q}\in\mathbb{R}^{n}\setminus\{\mathbf{p}\}$ is given by
@@
II-C3 Plane
Given two linearly independent vectors $\mathbf{p},\mathbf{q}\in\mathbb{R}^{n}$, the plane containing all vectors that can be expressed as a linear combination of $\mathbf{p}$ and $\mathbf{q}$ is given by
$\mathcal{P}(\mathbf{p},\mathbf{q}):=\{\mathbf{x}\in\mathbb{R}^{n}|\exists k_{1}\in\mathbb{R},\exists k_{2}\in\mathbb{R},\mathbf{x}=k_{1}\mathbf{p}+k_{2}\mathbf{q}\}$ (5)
II-C4 Hyperplane
@@
$\mathcal{H}(\mathbf{p},\mathbf{q}):=\{\mathbf{x}\in\mathbb{R}^{n}|\mathbf{q}^{\intercal}(\mathbf{x}-\mathbf{p})=0\}.$ (6)
The hyperplane divides the Euclidean space $\mathbb{R}^{n}$ into two half-spaces i.e., a closed positive half-space $\mathcal{H}_{\geq}(\mathbf{p},\mathbf{q})$ and a closed negative half-space $\mathcal{H}_{\leq}(\mathbf{p},\mathbf{q})$ which are obtained by substituting ‘$=$’ with $\geq$’ and $\leq$’ respectively, in the right-hand side of (6). We also use the notations $\mathcal{H}_{>}(\mathbf{p},\mathbf{q})$ and $\mathcal{H}_{<}(\mathbf{p},\mathbf{q})$ to denote the open positive and the open negative half-spaces such that $\mathcal{H}_{>}(\mathbf{p},\mathbf{q})=\mathcal{H}_{\geq}(\mathbf{p},\mathbf{q})\backslash\mathcal{H}(\mathbf{p},\mathbf{q})$ and $\mathcal{H}_{<}(\mathbf{p},\mathbf{q})=\mathcal{H}_{\leq}(\mathbf{p},\mathbf{q})\backslash\mathcal{H}(\mathbf{p},\mathbf{q})$.
II-C5 Cylinder
Given $\mathbf{p}\in\mathbb{R}^{n}$ and $w>0$, the cylinder with width $w$ and the line segment $\mathcal{L}_{s}(\mathbf{p},\mathbf{0})$ as its axis is denoted by
$\mathcal{C}_{\mathcal{L}}(\mathbf{p},w):=\mathcal{D}_{w}(\mathcal{L}_{s}(\mathbf{p},\mathbf{0}))\cap\mathcal{H}_{\geq}(\mathbf{p},-\mathbf{p})\cap\mathcal{H}_{\geq}(\mathbf{0},\mathbf{p})$
II-D Parallel projection operator
Given two orthonormal vectors $\mathbf{p},\mathbf{q}\in\mathbb{S}^{n-1}$, the parallel projection operator is defined as
$\mathbf{P}(\mathbf{p},\mathbf{q})=\mathbf{p}\mathbf{p}^{\top}+\mathbf{q}\mathbf{q}^{\top}.$
For any vector $\mathbf{s}\in\mathbb{R}^{n}$, the vector $\mathbf{s}^{\prime}$, where $\mathbf{s}^{\prime}=\mathbf{P}(\mathbf{p},\mathbf{q})\mathbf{s}$, is the projection of $\mathbf{s}$ onto the plane $\mathcal{P}(\mathbf{p},\mathbf{q})$. Additionally, it can be verified that $\mathbf{s}^{\top}\mathbf{s}^{\prime}\geq 0$, indicating that the angle between $\mathbf{s}$ and $\mathbf{s}^{\prime}$ is either acute or $90^{\circ}$.
II-E n-Dimensional rotation matrix
Consider two orthonormal vectors $\mathbf{p},\mathbf{q}\in\mathbb{S}^{n-1}$ and an angle $\theta\in[0,2\pi)$. The rotation matrix $R(\theta,\mathbf{p},\mathbf{q})$ is constructed as follows:
$R(\theta,\mathbf{p},\mathbf{q}):=\mathbf{I}_{n}+\sin(\theta)\mathbf{Sk}(\mathbf{p},\mathbf{q})+(1-\cos(\theta))(\mathbf{Sk}(\mathbf{p},\mathbf{q}))^{2},$
where the skew-symmetric matrix $\mathbf{Sk}(\mathbf{p},\mathbf{q})$ is defined by
$\mathbf{Sk}(\mathbf{p},\mathbf{q})=\mathbf{q}\mathbf{p}^{\top}-\mathbf{p}\mathbf{q}^{\top}.$
The operator $R(\theta,\mathbf{p},\mathbf{q})$ performs a rotation by the angle $\theta$ in the plane spanned by $\mathbf{p}$ and $\mathbf{q}$. Rotation angles are considered positive if they are performed from the vector $\mathbf{p}$ to the vector $\mathbf{q}$.
II-F Hybrid system framework
@@
[…]
A subset $\mathbb{T}\subset\mathbb{R}_{\geq}\times\mathbb{N}$ is a hybrid time domain if it is a union of a finite or infinite sequence of intervals $[t_{j},t_{j+1}]\times\{j\},$ where the last interval (if existent) is possibly of the form $[t_{j},T)$ with $T$ finite or $T=+\infty$. The ordering of points on each hybrid time domain is such that $(t,j)\preceq(t^{\prime},j^{\prime})$ if $t\leq t^{\prime}$ and $j\leq j^{\prime}$. A hybrid solution $\phi$ is maximal if it cannot be extended, and complete if its domain dom $\phi$ (which is a hybrid time domain) is unbounded.